OpenStudy (anonymous):

the quadratic equation ax2+bx= -1/2 has two equal roots. express a in terms of b

OpenStudy (anonymous):

ax^2 + bx + 1/2 = 0 delta = b^2 - 4a(1/2) = 0 -> b^2 - 2a = 0 => b^2 = 2a => x = -b /2a = -b/b^2 = -1/b

OpenStudy (anonymous):

a=b^2/2

OpenStudy (anonymous):

\( b^2=2a \)